Solve the following equation by using FOIL:
x^2-2x-35=0

Found 2 solutions by rapaljer, stanbon:
Answer by rapaljer(4671)

This one FACTORS!!

x^2 - 2x - 35 = 0?
(x-___)(x+___) = 0
(x - 7)(x + 5) = 0

Now, you have TWO small equations to solve:
x-7=0
x=7

x+5=0
x=-5

You may want to see my FREE website for additional explanation on this topic. The easiest way to find the website is to use the easy-to-remember and easy-to-spell link www.mathinlivingcolor.com. Near the bottom of this page is a link that takes you to my Homepage. On my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra" and look in Chapter 2 for the section on Factoring Trinomial and also the section on Solving Quadratic Equations by Factoring. I think you will really like the "Math in Living Color" pages that go with these sections. In addition, I have a FREE video on this topic that was made in my younger years before I retired. To see the video, look on my Homepage for the link "Rapalje Videos in Living Color." Choose Basic Algebra, and look for the video on "Factoring."

If you need to contact me, send me an Email at rapaljer@seminolestate.edu.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus

Answer by stanbon(75887)

Solve the following equation by using FOIL:
x^2-2x-35=0
---
Using inverse-FOIL you get:
x^2-7x+5x-35 = 0
x(x-7)+5(x-7) = 0
(x-7)(x+5) = 0
x = 7 or x = -5
======================
Cheers,
Stan H.
================